Which of these two solutions will have the highest boiling point: C12H22O11 0.50 mol / L or KCl 0.50 mol / L? Explain your reasoning + show steps
Answer #1

The expression of elevation in boiling point is given as:

\Delta T_b=i\times k_b\times m

where,

\Delta T_b = Elevation in boiling point

i = Van't Hoff factor

T_b = change in boiling point

k_b = boiling point constant

m = molality

For C12H22O11 0.50 mol / L solution

i= 1,

So molar concentration will be (1× 0.5)= 0.5 m

For KCl 0.50 mol / L solution

i= 2,

So molar concentration will be (2×0.5)= 1 m

As, the molal concentration of KCl 0.50 mol / L solution is the highest, so its boiling point will be the highest.